# A mixed Petrov–Galerkin Cosserat rod finite element formulation

### v2

@@ 3.4 Constraint deformations and static condensation
containing the same equations as ${\bm{K}}_{c}^{-1}{\bm{\lambda}}_{c}-{\bm{l}}_{c}({\bm{q}})=\mathbf{0}_{6pn_{\mathrm{el}}}$ but with the matrix $\bar{{\bm{K}}}_{c}^{-1}$ having full rank and thus being invertible. The nodal resultant contact forces and moments of ${\bm{\lambda}}_{c}$ that are not contained in $\bar{{\bm{\lambda}}}_{c}$ are then the Lagrange multipliers required to satisfy ${\bm{g}}_{c}({\bm{q}})=\mathbf{0}_{n_{\mathrm{g}}pn_{\mathrm{el}}}$. Thus, the constraints derived through this method are equivalent to constrained rod finite element models based on the Lagrange multiplier methods, see for example [32]. Furthermore, we can also use only the constraints from the mixed formulation and keep the unconstrained deformations in the displacement based formulation. This approach will be later used in the experiment presented in Section˜4.6, where we show the case of a non-convex strain energy density function together with constraints.
@@
Following the argumentation of [51, 45], the solution of the generalized coordinates $\bm{q}$} using the displacement-based formulation is equivalent to the solution of the generalized coordinates $\bm{q}$} using the mixed formulation, when in both cases reduced integration is applied. Since the resultant contact forces and moments are interpolated with polynomial degree p1p-1, and therefore having $p$ unknowns per component, this matches the number of quadrature points used in case of reduced integration. In the mixed formulation, the resultant contact forces and moments therefore fulfill at these quadrature points the constitutive equations $_{B}{\bm{n}}={\bm{C}}_{\bm{\gamma}}\,{\bm{\varepsilon}}_{\bm{\gamma}}$ and $_{B}{\bm{m}}={\bm{C}}_{\bm{\kappa}}\,{\bm{\varepsilon}}_{\bm{\kappa}}$. In the displacement-based formulation, the constitutive equations are fulfilled at all points and especially also at the quadrature points. Consequently, also the static condensed formulation has the same solution $\bm{q}$}, and it must hold therefore ${\bm{f}}^{\mathrm{int}}({\bm{q}})={\bm{W}}_{c}({\bm{q}}){\bm{K}}_{c}\,{\bm{l}}_{c}({\bm{q}})$ in case of reduced integration.
We want to highlight, that although the solution of the generalized coordinates $\bm{q}$} of the static condensed formulation and the mixed formulation are identical, the iterates taken by the solver are different. This is also in agreement with [1], where the mixed formulation naturally introduces the auxiliary variables leading to a much better convergence when the Newton–Raphson method is applied. Note also that, due to numerical computations, the iterations and solutions of the different formulations differ slightly.
@@ 3.5 SE3{SE}3 interpolation
Additionally, the SE3{SE}3 interpolation outlined in [34, 24] will be analyzed. The formulation is adapted to also use a quaternion parametrization for the nodal transformation matrices instead of rotation vectors. Therefore, the quaternion constraints ${\bm{g}}_{S}({\bm{q}})$ are also introduced. Since the SE3{SE}3 interpolation leads to element-wise constant strains, a mixed internal virtual work with element-wise constant resultant contact forces and moments is applied. The tuples $\bm{q}$}, $\delta{\bm{s}}$, ${\bm{\lambda}}_{c}$ and $\delta{\bm{\lambda}}_{c}$ are therefore identical to the ones used in the quaternion interpolation with polynomial degree $p=1$.
@@ 4 Numerical experiments
Interpolation
Quaternion, $p=1$ $\mathcal{Q}^{1}_{DB_{\mathrm{full}}}$ $\mathcal{Q}^{1}_{DB_{\mathrm{red}}}$ $\mathcal{Q}^{1}_{MX_{\mathrm{full}}}$ $\mathcal{Q}^{1}_{MX_{\mathrm{red}}}$
@@ 4.2 Helix experiment
This benchmark of the helix experiment was used in [48], [32] and [33]. In this work, it is used to demonstrate the computational efficiency of the mixed approach. A helical curve with $n=2$ coils along the ${\bm{e}}_{z}^{I}$ axis with height $h=50$ and radius $R_{0}=10$ is described with $\xi\in[0,1]$ by
@@
where $c=h/(2\pi R_{0}n)$ is the pitch of the helix. The length of the curve is $L=2\pi R_{0}n\sqrt{1+c^{2}}$. The considered rod is initially straight with length $L$ and has a circular cross-section with radius $r=L/(2\rho)$, where $\rho\in\{10^{1},10^{2},10^{3},10^{4}\}$ is the slenderness ratio. The stiffnesses are given by $k_{\mathrm{e}}=EA$, $k_{\mathrm{s}_{y}}=k_{\mathrm{s}_{z}}=GA$, $k_{\mathrm{t}}=2GI$ and $k_{\mathrm{b}_{y}}=k_{\mathrm{b}_{z}}=EI$ in terms of Young’s and shear moduli $E=1$ and $G=1/2$, the cross-section area $A=\pi r^{2}$ and the second moment of area $I=\pi r^{4}/4$. Note that by this choice, it holds $k_{\mathrm{t}}=k_{\mathrm{b}_{y}}=k_{\mathrm{b}_{z}}$. The rod is clamped at $\xi=0$, such that $_{I}{\bm{r}}_{OC}(0)=(0,-R_{0},0)$, $_{I}{\bm{e}}_{x}^{B}(0)=(1,0,c)/\sqrt{1+c^{2}}$, $_{I}{\bm{e}}_{y}^{B}(0)=(0,1,0)$ and $_{I}{\bm{e}}_{z}^{B}(0)={}_{I}{\bm{e}}_{x}^{B}(0)\times{}_{I}{\bm{e}}_{y}^{B}(0)$. To bend the rod into the shape from ˜29, a moment $_{B}{\bm{c}}_{1}=(ck_{\mathrm{t}},0,k_{\mathrm{b}_{z}})/\big(R_{0}(1+c^{2})\big)$ is applied at the free end at $\xi=1$. The analytical solution has constant resultant contact forces $_{B}{\bm{n}}(\xi)=\mathbf{0}$ and moments $_{B}{\bm{m}}(\xi)={}_{B}{\bm{c}}_{1}$, as discussed in [32]. All helix experiments were performed using $N=17$ nodes. In Figure˜4, the final deformed configuration of the slenderness ratios $\rho\in\{10^{1},10^{2}\}$ are shown.
@@
Figure˜5 shows the resultant contact forces and moments for the slenderness ratio $\rho=10^{1}$, obtained with 128 load increments for displacement-based formulations and with $1$ load increment for mixed formulations. The resultant contact forces and moments of both quaternion interpolations with $DB_{\mathrm{full}}$ indicate locking, which can be seen as the resultant contact moment is not “periodic” with respect to the centerline parameter. In fact, the deformed rod is not even close to the desired helical shape. The resultant contact forces and moments of both quaternion interpolations with $DB_{\mathrm{red}}$ show the typical fluctuation around the exact value, as already noted in the previous experiment. The resultant contact forces and moments for $MX_{\mathrm{full}}$ and $MX_{\mathrm{red}}$ are almost identical and represent in each case the expected solution. All formulations of the SE3{SE}3 rod are not only able to represent the correct resultant contact forces and moments, but also to correctly represent the helix shape, as discussed in [32].
@@ 4.3 Rod bent to a helical form
The next benchmark was proposed the first time in [36] to simulate large inhomogeneous deformations. Here it is used to validate the shown computational improvements from the previous section on another problem. In this experiment, an initially straight cantilever with length $L=10$ is considered. The stiffnesses are given by $k_{\mathrm{e}}=k_{\mathrm{s}_{y}}=k_{\mathrm{s}_{z}}=10^{4}$ and $k_{\mathrm{t}}=k_{\mathrm{b}_{y}}=k_{\mathrm{b}_{z}}=10^{2}$. On the cantilever’s free tip an external moment $_{B}{\bm{c}}_{1}={\bm{A}}_{IB}^{\!\mathrm{T}}(0,0,20\pi k_{\mathrm{b}_{y}}/L)$ and an external force $_{I}{\bm{b}}_{1}=(0,0,50)$ are applied, see Figure˜7. Note that both the moment and the force are constant w.r.t. the inertial $I$-system, resulting in a generalized external force which depends on the generalized coordinates $\bm{q}$}. As the stiffness parameters are not computed by geometric properties, the shown circular cross-section is only for visualization. Further, note that without the perturbation force $_{I}{\bm{b}}_{1}$, the external moment would roll-up the cantilever to 10 coils, where the resultant contact forces and moments would be constant. The tolerance of Newton’s solver was set to $\epsilon=10^{-8}$. We used for every rod $N=61$ nodes, leading to $n_{\mathrm{el}}=60$ elements for the $\mathcal{Q}^{1}$ and SE3{SE}3 formulations and $n_{\mathrm{el}}=30$ elements for the $\mathcal{Q}^{2}$ formulation. Because $DB_{\mathrm{full}}$ formulations show locking behavior, and for mixed formulations the reduced integration method leads to almost the same results as the full integration, we will only compare the $DB_{\mathrm{red}}$ and the $MX_{\mathrm{full}}$ formulations in this experiment.
@@ 4.4 Deployment of an elastic ring
In this benchmark experiment, a circular ring with rectangular cross-section is deployed. It has been proposed by [69] to analyze the buckling phenomenon of a ring when exposed to a moment. It is used as benchmark experiment by various authors [[, see]]DaCostaESilva2020,Greco2022,Greco2024,Romero2004,Meier2014,Smolenski1999,Borkovic2023. The centerline of the ring in the reference configuration has a circular shape with radius $R=20$. Point $\mathcal{A}$ is located at $\xi=0$, where the ring is clamped to the origin. To keep the ring closed in the simulation, the other end of the ring is also clamped. At point $\mathcal{C}$, which is located at $\xi=1/2$, an external moment $M$ along the ${\bm{e}}_{x}^{I}$ axis is applied, such that the rotated angle around this axis matches a prescribed angle $\theta$. We introduce the basis at point $\mathcal{C}$ by ${\bm{A}}_{I\mathcal{C}}={\bm{A}}_{IB}(\xi=1/2)$. Furthermore, point $\mathcal{B}$ [^2] In literature, the points are labeled as $A$, $B$ and $C$. To avoid confusion with the centerline curve $_{I}{\bm{r}}_{OC}$ and the transformation matrix describing the cross-section orientation ${\bm{A}}_{IB}$, we label them $\mathcal{A}$, $\mathcal{B}$ and $\mathcal{C}$, respectively. is located at $\xi=1/4$ and is used to investigate the deformation of the ring. The placement of point $\mathcal{B}$ is introduced as $_{I}{\bm{r}}_{O\mathcal{B}}={}_{I}{\bm{r}}_{OC}(\xi=1/4)$. The setup of the experiment is shown in Figure˜10. The cross-section of the ring has height $h=1$ and width in radial direction $w=1/3$. With Young’s modulus $E=2.1\cdot 10^{7}$ and Poisson’s ratio $\nu=0.3$, the shear modulus is given by $G=E/(2(1+\nu))$. The stiffnesses are given by $k_{\mathrm{e}}=EA$, $k_{\mathrm{s}_{y}}=k_{\mathrm{s}_{z}}=GA$, $k_{\mathrm{t}}=G\cdot 9.753\cdot 10^{-3}$, $k_{\mathrm{b}_{y}}=EI_{y}$ and $k_{\mathrm{b}_{z}}=EI_{z}$ with $A=hw$, $I_{y}=wh^{3}/12$ and $I_{z}=hw^{3}/12$. Note that the value for the torsional stiffness $k_{\mathrm{t}}$ is not given by the geometric value of the cross-section, instead a custom value, as done by [31, 30], is taken.
@@
The ring is discretized with $n_{\mathrm{el}}=20$ elements of the $\mathcal{Q}^{2}_{MX_{\mathrm{full}}}$ formulation. Since the number of elements is even, point $\mathcal{C}$ is located at the boundary of adjacent elements so that the requirement on external loads is fulfilled as described after ˜23. The tolerance for Newton’s solver is given by $\epsilon=10^{-6}$, and the angle of rotation in point $\mathcal{C}$ is increased within 120 increments from $\theta=0$ to $\theta=4\pi$. Figure˜10 shows the required moment $M$ to enforce the rotated angle $\theta$ at point $\mathcal{C}$. The configurations for different rotated angles are shown in Figure˜11. The components of the displacement of point $\mathcal{B}$ are shown in Figure˜12, where the components are introduced as $(\Delta r_{x},\Delta r_{y},\Delta r_{z})={}_{I}{\bm{r}}_{O\mathcal{B}}-{}_{I}{\bm{r}}_{O\mathcal{B}}^{0}$. Due to symmetry, one could also just simulate one half of the ring. This divides also the required moment $M$ by two, see for instance [18, 30]. The results obtained using the herein proposed rod formulation show good agreement with results from literature.

### v3

@@ 3.4 Constraint deformations and static condensation
containing the same equations as ${\bm{K}}_{c}^{-1}{\bm{\lambda}}_{c}-{\bm{l}}_{c}({\bm{q}})=\mathbf{0}_{6pn_{\mathrm{el}}}$ but with the matrix $\bar{{\bm{K}}}_{c}^{-1}$ having full rank and thus being invertible. The nodal resultant contact forces and moments of ${\bm{\lambda}}_{c}$ that are not contained in $\bar{{\bm{\lambda}}}_{c}$ are then the Lagrange multipliers required to satisfy ${\bm{g}}_{c}({\bm{q}})=\mathbf{0}_{n_{\mathrm{g}}pn_{\mathrm{el}}}$. Thus, the constraints derived through this method are equivalent to constrained rod finite element models based on the Lagrange multiplier methods, see for example [undefae]. Furthermore, we can also use only the constraints from the mixed formulation and keep the unconstrained deformations in the displacement based formulation. This approach will be later used in the experiment presented in Section˜4.6, where we show the case of a non-convex strain energy density function together with constraints.
@@
Following the argumentation of [undefax, undefar], the solution of the generalized coordinates $\bm{q}$} using the displacement-based formulation is equivalent to the solution of the generalized coordinates $\bm{q}$} using the mixed formulation, when in both cases reduced integration is applied. Since the resultant contact forces and moments are interpolated with polynomial degree p1p-1, and therefore having $p$ unknowns per component, this matches the number of quadrature points used in case of reduced integration. In the mixed formulation, the resultant contact forces and moments therefore fulfill at these quadrature points the constitutive equations $_{B}{\bm{n}}={\bm{C}}_{\bm{\gamma}}\,{\bm{\varepsilon}}_{\bm{\gamma}}$ and $_{B}{\bm{m}}={\bm{C}}_{\bm{\kappa}}\,{\bm{\varepsilon}}_{\bm{\kappa}}$. In the displacement-based formulation, the constitutive equations are fulfilled at all points and especially also at the quadrature points. Consequently, also the static condensed formulation has the same solution $\bm{q}$}, and it must hold therefore ${\bm{f}}^{\mathrm{int}}({\bm{q}})={\bm{W}}_{c}({\bm{q}}){\bm{K}}_{c}\,{\bm{l}}_{c}({\bm{q}})$ in case of reduced integration.
We want to highlight, that although the solution of the generalized coordinates $\bm{q}$} of the static condensed formulation and the mixed formulation are identical, the iterates taken by the solver are different. This is also in agreement with [undef], where the mixed formulation naturally introduces the auxiliary variables leading to a much better convergence when the Newton–Raphson method is applied. Note also that, due to numerical computations, the iterations and solutions of the different formulations differ slightly.
@@ 3.5 SE3{SE}3 interpolation
Additionally, the SE3{SE}3 interpolation outlined in [undefag, undefw] will be analyzed. The formulation is adapted to also use a quaternion parametrization for the nodal transformation matrices instead of rotation vectors. Therefore, the quaternion constraints ${\bm{g}}_{S}({\bm{q}})$ are also introduced. Since the SE3{SE}3 interpolation leads to element-wise constant strains, a mixed internal virtual work with element-wise constant resultant contact forces and moments is applied. The tuples $\bm{q}$}, $\delta{\bm{s}}$, ${\bm{\lambda}}_{c}$ and $\delta{\bm{\lambda}}_{c}$ are therefore identical to the ones used in the quaternion interpolation with polynomial degree $p=1$.
@@ 4 Numerical experiments
Interpolation Quaternion, $p=1$ $\mathcal{Q}^{1}_{DB_{\mathrm{full}}}$ $\mathcal{Q}^{1}_{DB_{\mathrm{red}}}$ $\mathcal{Q}^{1}_{MX_{\mathrm{full}}}$ $\mathcal{Q}^{1}_{MX_{\mathrm{red}}}$
@@ 4.2 Helix experiment
This benchmark of the helix experiment was used in [undefau], [undefae] and [undefaf]. In this work, it is used to demonstrate the computational efficiency of the mixed approach. A helical curve with $n=2$ coils along the ${\bm{e}}_{z}^{I}$ axis with height $h=50$ and radius $R_{0}=10$ is described with $\xi\in[0,1]$ by
@@
where $c=h/(2\pi R_{0}n)$ is the pitch of the helix. The length of the curve is $L=2\pi R_{0}n\sqrt{1+c^{2}}$. The considered rod is initially straight with length $L$ and has a circular cross-section with radius $r=L/(2\rho)$, where $\rho\in\{10^{1},10^{2},10^{3},10^{4}\}$ is the slenderness ratio. The stiffnesses are given by $k_{\mathrm{e}}=EA$, $k_{\mathrm{s}_{y}}=k_{\mathrm{s}_{z}}=GA$, $k_{\mathrm{t}}=2GI$ and $k_{\mathrm{b}_{y}}=k_{\mathrm{b}_{z}}=EI$ in terms of Young’s and shear moduli $E=1$ and $G=1/2$, the cross-section area $A=\pi r^{2}$ and the second moment of area $I=\pi r^{4}/4$. Note that by this choice, it holds $k_{\mathrm{t}}=k_{\mathrm{b}_{y}}=k_{\mathrm{b}_{z}}$. The rod is clamped at $\xi=0$, such that $_{I}{\bm{r}}_{OC}(0)=(0,-R_{0},0)$, $_{I}{\bm{e}}_{x}^{B}(0)=(1,0,c)/\sqrt{1+c^{2}}$, $_{I}{\bm{e}}_{y}^{B}(0)=(0,1,0)$ and $_{I}{\bm{e}}_{z}^{B}(0)={}_{I}{\bm{e}}_{x}^{B}(0)\times{}_{I}{\bm{e}}_{y}^{B}(0)$. To bend the rod into the shape from ˜29, a moment $_{B}{\bm{c}}_{1}=(ck_{\mathrm{t}},0,k_{\mathrm{b}_{z}})/\big(R_{0}(1+c^{2})\big)$ is applied at the free end at $\xi=1$. The analytical solution has constant resultant contact forces $_{B}{\bm{n}}(\xi)=\mathbf{0}$ and moments $_{B}{\bm{m}}(\xi)={}_{B}{\bm{c}}_{1}$, as discussed in [undefae]. All helix experiments were performed using $N=17$ nodes. In Figure˜4, the final deformed configuration of the slenderness ratios $\rho\in\{10^{1},10^{2}\}$ are shown.
@@
Figure˜5 shows the resultant contact forces and moments for the slenderness ratio $\rho=10^{1}$, obtained with 128 load increments for displacement-based formulations and with $1$ load increment for mixed formulations. The resultant contact forces and moments of both quaternion interpolations with $DB_{\mathrm{full}}$ indicate locking, which can be seen as the resultant contact moment is not “periodic” with respect to the centerline parameter. In fact, the deformed rod is not even close to the desired helical shape. The resultant contact forces and moments of both quaternion interpolations with $DB_{\mathrm{red}}$ show the typical fluctuation around the exact value, as already noted in the previous experiment. The resultant contact forces and moments for $MX_{\mathrm{full}}$ and $MX_{\mathrm{red}}$ are almost identical and represent in each case the expected solution. All formulations of the SE3{SE}3 rod are not only able to represent the correct resultant contact forces and moments, but also to correctly represent the helix shape, as discussed in [undefae].
@@ 4.3 Rod bent to a helical form
The next benchmark was proposed the first time in [undefai] to simulate large inhomogeneous deformations. Here it is used to validate the shown computational improvements from the previous section on another problem. In this experiment, an initially straight cantilever with length $L=10$ is considered. The stiffnesses are given by $k_{\mathrm{e}}=k_{\mathrm{s}_{y}}=k_{\mathrm{s}_{z}}=10^{4}$ and $k_{\mathrm{t}}=k_{\mathrm{b}_{y}}=k_{\mathrm{b}_{z}}=10^{2}$. On the cantilever’s free tip an external moment $_{B}{\bm{c}}_{1}={\bm{A}}_{IB}^{\!\mathrm{T}}(0,0,20\pi k_{\mathrm{b}_{y}}/L)$ and an external force $_{I}{\bm{b}}_{1}=(0,0,50)$ are applied, see Figure˜7. Note that both the moment and the force are constant w.r.t. the inertial $I$-system, resulting in a generalized external force which depends on the generalized coordinates $\bm{q}$}. As the stiffness parameters are not computed by geometric properties, the shown circular cross-section is only for visualization. Further, note that without the perturbation force $_{I}{\bm{b}}_{1}$, the external moment would roll-up the cantilever to 10 coils, where the resultant contact forces and moments would be constant. The tolerance of Newton’s solver was set to $\epsilon=10^{-8}$. We used for every rod $N=61$ nodes, leading to $n_{\mathrm{el}}=60$ elements for the $\mathcal{Q}^{1}$ and SE3{SE}3 formulations and $n_{\mathrm{el}}=30$ elements for the $\mathcal{Q}^{2}$ formulation. Because $DB_{\mathrm{full}}$ formulations show locking behavior, and for mixed formulations the reduced integration method leads to almost the same results as the full integration, we will only compare the $DB_{\mathrm{red}}$ and the $MX_{\mathrm{full}}$ formulations in this experiment.
@@ 4.4 Deployment of an elastic ring
In this benchmark experiment, a circular ring with rectangular cross-section is deployed. It has been proposed by [undefaap] to analyze the buckling phenomenon of a ring when exposed to a moment. It is used as benchmark experiment by various authors [[, see]]DaCostaESilva2020,Greco2022,Greco2024,Romero2004,Meier2014,Smolenski1999,Borkovic2023. The centerline of the ring in the reference configuration has a circular shape with radius $R=20$. Point $\mathcal{A}$ is located at $\xi=0$, where the ring is clamped to the origin. To keep the ring closed in the simulation, the other end of the ring is also clamped. At point $\mathcal{C}$, which is located at $\xi=1/2$, an external moment $M$ along the ${\bm{e}}_{x}^{I}$ axis is applied, such that the rotated angle around this axis matches a prescribed angle $\theta$. We introduce the basis at point $\mathcal{C}$ by ${\bm{A}}_{I\mathcal{C}}={\bm{A}}_{IB}(\xi=1/2)$. Furthermore, point $\mathcal{B}$ [^2] In literature, the points are labeled as $A$, $B$ and $C$. To avoid confusion with the centerline curve $_{I}{\bm{r}}_{OC}$ and the transformation matrix describing the cross-section orientation ${\bm{A}}_{IB}$, we label them $\mathcal{A}$, $\mathcal{B}$ and $\mathcal{C}$, respectively. is located at $\xi=1/4$ and is used to investigate the deformation of the ring. The placement of point $\mathcal{B}$ is introduced as $_{I}{\bm{r}}_{O\mathcal{B}}={}_{I}{\bm{r}}_{OC}(\xi=1/4)$. The setup of the experiment is shown in Figure˜10. The cross-section of the ring has height $h=1$ and width in radial direction $w=1/3$. With Young’s modulus $E=2.1\cdot 10^{7}$ and Poisson’s ratio $\nu=0.3$, the shear modulus is given by $G=E/(2(1+\nu))$. The stiffnesses are given by $k_{\mathrm{e}}=EA$, $k_{\mathrm{s}_{y}}=k_{\mathrm{s}_{z}}=GA$, $k_{\mathrm{t}}=G\cdot 9.753\cdot 10^{-3}$, $k_{\mathrm{b}_{y}}=EI_{y}$ and $k_{\mathrm{b}_{z}}=EI_{z}$ with $A=hw$, $I_{y}=wh^{3}/12$ and $I_{z}=hw^{3}/12$. Note that the value for the torsional stiffness $k_{\mathrm{t}}$ is not given by the geometric value of the cross-section, instead a custom value, as done by [undefad, undefac], is taken.
@@
The ring is discretized with $n_{\mathrm{el}}=20$ elements of the $\mathcal{Q}^{2}_{MX_{\mathrm{full}}}$ formulation. Since the number of elements is even, point $\mathcal{C}$ is located at the boundary of adjacent elements so that the requirement on external loads is fulfilled as described after ˜23. The tolerance for Newton’s solver is given by $\epsilon=10^{-6}$, and the angle of rotation in point $\mathcal{C}$ is increased within 120 increments from $\theta=0$ to $\theta=4\pi$. Figure˜10 shows the required moment $M$ to enforce the rotated angle $\theta$ at point $\mathcal{C}$. The configurations for different rotated angles are shown in Figure˜11. The components of the displacement of point $\mathcal{B}$ are shown in Figure˜12, where the components are introduced as $(\Delta r_{x},\Delta r_{y},\Delta r_{z})={}_{I}{\bm{r}}_{O\mathcal{B}}-{}_{I}{\bm{r}}_{O\mathcal{B}}^{0}$. Due to symmetry, one could also just simulate one half of the ring. This divides also the required moment $M$ by two, see for instance [undefq, undefac]. The results obtained using the herein proposed rod formulation show good agreement with results from literature.